OpenStudy (anonymous):

Given that n(A)=15, n(B)=12, and n(A ∪ B)=17, what is n(A ∩ B)?

OpenStudy (zarkon):

\[n(A\cup B)=n(A)+n(B)-n(A\cap B)\] plug in your numbers and solve

OpenStudy (anonymous):

even for the n(A∪B) and n(A∩B)

OpenStudy (zarkon):

plug in the numbers you have... \[n(A\cup B)=n(A)+n(B)-n(A\cap B)\] \[17=15+12-n(A\cap B)\] now solve for the uinknown.

OpenStudy (zarkon):

*unknown

OpenStudy (anonymous):

So the answer would be 10?

OpenStudy (zarkon):

yes